# The Surprising Effectiveness of Test-Time Training for Abstract Reasoning

### v1

@@ 2.4 Experimental Setup
We use an 8B parameter language model from the Llama-3 models, and 1B, 3B from Llama-3.2 models (Dubey et al., 2024). We use Low-Rank Adaptation (LoRA) (Hu et al., 2021) for parameter-efficient test-time training. For each task $d$, we initialize a separate set of LoRA parameters that are trained on the dataset $\mathcal{D}_{\textrm{TTT}}$. The LoRA rank is set to 128, and adaptations are applied to MLP, attention, and output layers. We train models with AdamW optimizer (Loshchilov and Hutter, 2019) with 2 epochs with batch sizes of 2.
Data & Formatting
@@
For efficient evaluation purposes, we randomly pick 80 balanced ARC tasks from ARC validation set, includes 20 easy, 20 medium, 20 hard, 20 expert tasks according to the classification in LeGris et al. (2024a) (see Table˜2 for this task list). We will use this subset of ARC tasks throughout the paper, except our final results given in for the full validation set (Table˜1). We limit $\mathcal{D}_{\textrm{TTT}}$ to have maximum of 250 examples per task for efficiency reasons. With that, the whole TTT and inference process takes approximately 12 hours for 100 randomly sampled validation tasks when using an NVIDIA-A100 GPU. Section˜B.2 provides additional details on the hyper-parameters. Input grids are converted to text using numpy’s default array printing format as shown in Fig.˜8.
In the following sections, we investigate the key factors that contribute to successful abstract reasoning with language models. Our analysis covers the impact of fine-tuning data $\mathcal{D}_{\textrm{FT}}$ data, TTT data $\mathcal{D}_{\textrm{TTT}}$, training objectives, inference procedures, and model size, providing insights into effective strategy for deploying test-time training.
3 What Dataset and Loss During TTT?
3.1 Data Generation
Given a task, we take the set of training input-output pairs $(x^{\textrm{train}}_{k},y^{\textrm{train}}_{k})_{k=1}^{K}$ and turn them into an augmented set of test-time-training tasks $\mathcal{D}_{\text{TTT}}$. We obtain $\mathcal{D}_{\text{TTT}}$ using a two-step process: First, we create a set of leave-one-out in-context learning tasks from the given training input-output pairs. Second, we use invertible rule-based transformations on this set to obtain an augmented dataset. This process is summarized in Fig.˜2.
Step 1 - Leave-one-out Tasks: By excluding the $j^{\textrm{th}}$ example pair from the training examples, we can create the following synthetic task:
$d_{j}^{\textrm{ICL}}=(\underbrace{(x_{k},y_{k})_{k\in\{1,...,K\}\setminus\{j\}}}_{\textrm{synthetic training examples}},\underbrace{{x_{j},\quad\quad y_{j}}}_{\textrm{synthetic test example}})\quad\text{where }j\in[1,K]$ (2)
where $d_{j}$ synthetic training task with the $j$-th example pair treated as the test case. We can generate $n$ different tasks, each containing n1n-1 example pairs. We further include two randomly permuted version of $d_{j}$ where we permute the order of the training examples.
Step 2 - Rule-based transformations: Consider an invertible transformation $t$ such that $t^{-1}(t(x))=x$. For every task obtained in step 1, we can use $t$ to generate a new augmented task $t(d_{j}^{\textrm{ICL}})$, where $t$ is applied to each individual grid in the task.
We choose simple transformations that preserve the fundamental relationships while introducing controlled variations such as rotation, flips, color permutation, example permutaton, size scaling, etc. The list and the description of these transformations are provided in Table˜3. Finally, we obtain
$\mathcal{D}_{\textrm{TTT-ICL}}=\{t(d_{j}^{\textrm{ICL}})\}\quad\text{for all $t,j$ pairs.}$ (3)
Baseline: End-to-End Learning Tasks
For comparison to the “test-time in-context learning” approach described above, we also evaluate an “test-time end-to-end learning” approach. We create a supervised dataset directly from the example demonstrations by treating each input-output pair as an independent training instance. Unlike the in-context learning setup, no context is used for prediction:
$d_{j}^{\textrm{E2E}}=(x_{j},y_{j})\quad\text{where }j\in[1,K]$ (4)
Note that this would be equivalent to leave-(n1n-1)-out task set in ICL setting as no training examples are provided as context. Similar to ICL case, we can apply rule-based transformations to augment the dataset:
$\mathcal{D}_{\textrm{TTT-E2E}}=\{t(d_{j}^{\textrm{E2E}})\}\quad\text{for all $t,j$ pairs.}$ (5)
This approach is computationally more efficient as it directly learns the input-output mapping without the overhead of managing demonstration context i.e. the few-shot prompt.
3.2 Optimization Objective
During test-time training, we optimize a set of task-specific parameters using LoRA (low-rank adaptation; Hu et al. (2021)) while keeping most of the base model frozen. This approach allows computationally efficient adaptation while maintaining the model’s general capabilities.
Training Objective:
Given a task’s test-time training dataset $\mathcal{D}^{i}_{\textrm{TTT}}$, we minimize the standard language modeling loss on both the demonstrations and test outputs:
$\mathcal{L}_{i}(\mathcal{D}^{i}_{\textrm{TTT}};{\bm{\theta}}_{\textrm{i}})=\sum_{d\in\mathcal{D}^{i}_{\textrm{TTT}}}\left(\sum_{n=2}^{K}\mathcal{L}_{\textrm{LM}}(y_{n}|x_{1},y_{1},...,x_{n};{\bm{\theta}}_{\textrm{i}})+\mathcal{L}_{\textrm{LM}}(y_{\textrm{test}}|x_{1},y_{1},...,x_{K},y_{K},x_{\textrm{test}};{\bm{\theta}}_{\textrm{i}})\right)$ (6)
where $\mathcal{L}_{\textrm{LM}}$ is the standard cross-entropy loss for language modeling. Note that we include loss terms for demonstrations starting from the second example ($n=2$). By doing so, we encourage the model to start reasoning about the transformation pattern from the second demonstration pair itself.
Task-Specific Parameters:
Instead of learning a single LoRA adapter for all tasks in the test set, we learn an individual task-specific LoRA adapter for each task. That is, we obtain $N$ different LoRA adapters, where $N$ is the number of test tasks.
3.3 Results
Figure 3: Accuracy of different data and optimization ablations in TTT: Our data ablation studies reveal that the ICL data format is crucial for effective TTT, and that applying transformations to augment the TTT dataset notably enhances performance. In optimization ablations, learning task-specific adapters significantly outperforms using a single adapter. Additionally, taking a loss on the in-context demonstrations provides a minor performance boost, while using quantized LoRA results in only a slight performance decrease drop. Full discussion in Section 3.
We compare the main implementation of our method to the following ablations:
  1. 1.
    FT (No TTT): The vanilla baseline where TTT is ablated and the fine-tuned model is used instead.
  2. 2.
    No Transformations: No transformation-based data augmentation. That is, data from step 2 of the data generation pipeline described in Section 3.1 is not included in the test-time training dataset.
  3. 3.
    End-to-End (E2E) Data: Instead of the standard in-context task setup, we use the end-to-end task formulation, as described in Section 3.1.
  4. 4.
    Shared TTT: In contrast to learning a task-specific LoRA adapter, a single LoRA adapter is learned using an aggregated dataset of all tasks.
  5. 5.
    No Demonstration Loss: No loss is taken on the demonstrations in the training outputs of the data. That is, the TTT loss is simply:
    $\mathcal{L}_{i}(\mathcal{D}^{i}_{\textrm{TTT}};{\bm{\theta}}_{\textrm{i}})=\sum_{d\in\mathcal{D}^{i}_{\textrm{TTT}}}\left(\mathcal{L}_{\textrm{LM}}\left(y_{\textrm{test}}|x_{1},y_{1},...,x_{K},y_{K},x_{\textrm{test}};{\bm{\theta}}_{\textrm{i}}\right)\right)$ (7)
  6. 6.
    QLoRA: Rather than full-precision base model updates, quantized LoRA adapters (Dettmers et al., 2024) are learned for each task, which is the alternative for LoRA considered for memory efficiency.
[…]
Next, we ablate multiple components of TTT optimization to analyze their contribution to the performance. Learning a single LoRA adapter across all tasks reduces performance on 7 tasks ($\bf 24\%$ decrease). This is expected as learning a dedicated adapter allows more parameters to train per task. Second, the decision that we made by taking loss on the output demonstrations marginally improves the performance ($26\to 29$), as we believe that this forces the model to reason about the transformation while processing the demonstrations. Finally, we observe that using quantized LoRA (QLoRA) only leads to a marginal drop in performance ($29\to 26$) — in memory-bottlenecked scenarios using QLoRA may be viable.
4 What Inference Strategy After TTT?
@@
Figure 4: Augmented inference and hierarchical voting (Section˜4): We use leave-one-out tasks and invertible geometric transformations to obtain multiple equivalent versions of the task for augmented inference. Predictions from these versions are aggregated with a hierarchical voting strategy: first, voting is performed within each transformation, and then the top candidates from each transformation undergo global voting to yield the top two predictions.
4.1 Augmented Inference
For a given task with training examples ${(x_{k},y_{k})}_{k=1}^{K}$ and test input $x_{\textrm{test}}$, we use invertible geometric transformations to produce equivalent transformed versions of the task, as shown in Fig.˜3. Let $\mathcal{T}$ be some set set of invertible geometric transformations (e.g., rotations and reflections). For each transformation $t\in\mathcal{T}$, we apply $t$ to all training demonstrations and the test input and run our model with these transformed inputs. We then apply the inverse transformation to obtain the final prediction for that transformation.
[…]
We further augment our predictions by permuting the order of training examples. For each transformation $g$, we sample $n=2$ different permutations of the demonstration sequence, resulting in $n\cdot|\mathcal{T}|$ total predictions per task. This is to mitigate any bias in the model’s processing of the demonstration sequence. Bober-Irizar and Banerjee (2024) also find transpose and rotation is helpful to produce extra prediction candidates.
4.2 Ensembling Predictions (Voting Strategy)
  1. 1.
  2. 2.
4.3 Results
Figure 5: Accuracy of different invertible transformations and voting schema: Our analysis shows that while individual transformations generally perform at a modest level and are comparable to one another, aggregating across them through voting yields substantial improvements. Notably, a hierarchical voting strategy with two voting stages surpasses a flat voting approach. Our hierarchical method approaches oracle-level performance, demonstrating its effectiveness in accurately selecting the correct answer when present. Full discussion in Section 4.3.
To analyze the impact of augmented inference and voting, we run the following ablations:
  1. 1.
    Vanilla: This baseline follows a standard inference approach without any augmented inference or voting. It generates 2 predictions from the model for $2$ permutations of the task. This setup serves as a reference point to assess the benefits of our augmented inference and voting strategy.
  2. 2.
    Transformed Inference (Rotate/Transpose/Flip): Measures performance when predictions are generated solely from a specific transformed version of the task shown in Fig.˜5. This assesses the individual effectiveness of each transformation applied in isolation. Note that Vanilla can also be considered a part of this category, with the transformation being the identity function .
  3. 3.
    Hierarchical Voting: Our full pipeline, which includes both augmented inference and voting.
  4. 4.
    Flattened Voting: Instead of using a hierarchical voting strategy, we perform a single voting round on the full set of $n\cdot|\mathcal{T}|$ predictions to identify the $2$ most frequent predictions.
  5. 5.
    Oracle: The oracle selects the correct answer if it exists in the set of $n\cdot|\mathcal{T}|$ predictions. The oracle provides an upper-bound on the best performance possible if the voting procedure was perfect.
The results are summarized in Figure 5. As shown in the figure, the individual performance of specific transformed versions is generally poor, with the transpose transformation yielding the worst accuracy. However, aggregating across these transformations through voting procedures leads to significant improvements. This suggests that some tasks may be easier to solve in their transformed versions, and that using self-consistency (voting) for aggregation is generally beneficial, a finding also observed in prior work. Additionally, while the flattened voting procedure improves accuracy, our hierarchical voting procedure outperforms it. In fact, our hierarchical procedure is comparable to the oracle, indicating that hierarchical aggregation effectively selects the correct answer (when it exists) with high accuracy.
5 What Fine-Tuning Before TTT?
Figure 6: LLM based synthetic tasks generation: Given some seed task descriptions and task generator functions in Python, we generate more generator functions to produce novel tasks. We use three different approaches: (1) few-shot prompting with only generators, (2) few-shot prompting with generators and task descriptions, (3) two-stage approach: first generate free form descriptions, then condition on them to generate more generators (shown in Fig.˜9).
5.1 Preparing Fine-tuning Data
Hodel (2024) provides domain-specific language (DSL), ReARC, as well as the transformation $f_{i}$ that solves the task-$i$, and the data generation function $g_{i}$ that are implemented in this DSL for each training task in the $\mathcal{D_{\textrm{ARC}}^{\textrm{train}}}$ dataset. These functions enable sampling of new input-output pairs that maintains the same underlying transformation principle:
$d=(x,y)\sim\textrm{eval}(g_{i})$ (10)
where $d$ represents a newly generated input-output pair that can be solved using the same transformation function $f_{i}$ as the original task-ii [^2] We can verify the generated examples by asserting $f_{i}(x)=y$..
(a) Using Existing Generators
The generator functions $g$s in ReARC already provide an effective data augmentation tool by producing different instantiations of same tasks. We generate extra samples from these training tasks by running these codes many times and randomly splitting these new examples ($d\sim\textrm{eval}(g_{i})$) to set of train and test examples. These augmented examples are already provided with their DSL release.
(b) Few-shot Prompting an LLM
Additionally, we used several approaches to generate novel tasks using an LM (in our case an ensemble of GPT4 and GPT4-o).
$g^{\prime}\sim\textrm{LM}(g_{1},g_{2},\dots,g_{m})$ (11)
where $g^{\prime}$ is a new generator function and $g_{1},\dots,g_{m}$} are existing generator functions (shown in Fig.˜6)s. We sample different $m$ examples by uniformly from existing training set. We repeat this process multiple times to get a good amount of tasks.
where $s_{i}$ represents the description of task $i$.
To get the task descriptions, we manually created seed descriptions for 10 training tasks. These seed descriptions were then used to generate descriptions for the training and validation tasks through few-shot prompting. To increase diversity of tasks we use task descriptions with hierarchical fields (category, summary, and description). The process of getting these descriptions provided in the Section˜D.1.
Instead of jointly generating task descriptions and function generations, we additionally deployed a two-stage approach described as following:
$s^{\prime}$ $\sim\textrm{LM}(s_{1},s_{2},\dots s_{m})$ (13)
This approach first generates a task description $s^{\prime}$ and then conditions the generator creation on both existing task pairs and the new description. In total we collected 6426 generators with these LLM based approaches. We provide qualitative samples from these LM generated tasks in Fig.˜11
(c) Geometric Transformations
  • Input grids only: $(x,y)\rightarrow(t(x),y)$
  • Output grids only: $(x,y)\rightarrow(x,t(y))$
  • Both input and output: $(x,y)\rightarrow(t(x),t(y))$
The complete specification of transformations and their application details are provided in Section˜B.1. These transformations are applied randomly to variants of tasks with 30% of the time.
5.2 Results
We perform full fine-tuning 1B, 3B Llama 3.2 instruction-tuned, and 8B Llama 3 instruction-tuned using augmented data. The format and training objective is same as the ones described for TTT in Section˜2.4. Hyper-parameter details are given in Section˜B.2. We do the following ablations for augmented data:
  1. 1.
  2. 2.
    All: We use all methods described in Section 5.1, including ReARC, rule-based augmentation, and LM generation.
  3. 3.
    No-Geom: We remove geometric transformations from all tasks.
  4. 4.
Figure 7: Left: Accuracy when fine-tuning with different data sources. While all fine-tuned models perform similarly, their performance after TTT shows considerable variance. As expected, removing geometric transformations from the fine-tuning data reduces performance compared to the model trained on the full dataset. Surprisingly, excluding LM-generated data from fine-tuning actually outperforms the model trained on all data. Right: Performance results across different model sizes. As expected, performance of the base fine-tuned model improves with increasing model size, aligning with current scaling law trends. However, the scaling behavior after TTT is less clear. For instance, the final performance of the 1B and 3B models is identical after TTT. Full discussion in Section 5.2.
How does FT data affect TTT?
We compare models using different fine-tuning data in Fig.˜7. We find that the model trained on ReARC with rule-based augmentation achieves the strongest performance. Surprisingly, including LM-generated tasks hurts performance by 5%, indicating that current LM-based task generation methods may need more sophisticated filtering mechanisms as used in Li et al. (2024) (see their results in Table˜1). Finally, we find that FT performance shows little correlation with TTT performance.
Model Size and Scaling in TTT
@@ 5.2 Results
We show results using different model sizes in Fig.˜7. Increasing the model size consistently improves FT performance, with the 8B model achieving the highest accuracy of 36%. We also observe that TTT effectively closes the performance gap for smaller models, with the 1B and 3B models achieving similar accuracy after TTT.
6 ARC Benchmark and Comparison to Other Systems
Table 1: Scores of different systems on the ARC validation set: Our TTT pipeline improves base models consistently. We achieve 47.1% accuracy when applied to our fine-tuned model, 53% when applied to BARC model from Li et al. (2024), achieving state-of-the-art on pure LM based approaches. We ensemble our method with program synthesis based models, where we achieve (61.87561.875%) state-of-the-art performance comparable to average human performance (60.2%).
Program Synthesizer Fine-tuned LM TTT Method Score (pass@2)
X Ours X 18.2518.25%
X Ours Ours 47.12547.125%
X BARC Ours 5353%
BARC Ours Ours 58.558.5%
BARC BARC Ours 61.87561.875%
Avg. Human 60.260.2%
Best Human 97.897.8%
BARC (ensemble) 54.37554.375%
BARC (no synthesizer) 39.2539.25%
Claude - Few-shot prompting 2121%
GPT-4.0 - Few-shot prompting 99%
@@ X BARC Ours 5353%
Following our development experiments on 80 tasks, we present comprehensive results on the full ARC public evaluation set, comparing our system against existing approaches. Our analysis focuses on three key aspects: the impact of our TTT methodology, the benefits of combining our approach with existing methods and the differences between fully neural and program synthesis methods.
Impact of Test Time Training
@@
We applied to our TTT and inference procedure (explained in Section˜3 and Section˜4) to our base fine-tuned models (fine-tuned 8B model without any LM data in Section˜5). TTT improves accuracy from 39.2539.25% to 47.12547.125%, surpassing existing end-to-end neural model results.
@@
A concurrent work by Li et al. (2024) introduced BARC, achieving 54.37554.375%accuracy by combining neural and program synthesis approaches—previously the highest publicly available result. While their fully neural approach shares similarities with our system, our TTT and inference pipeline has several additional components that boost performance. In particular, our test-time-training includes per-task LoRA and a larger set of augmentations, while our prediction pipeline includes an augmented inference under invertible transformations and a hierarchical self-consistency voting scheme. To validate our improvements, we applied our TTT pipeline to BARC’s fully neural model, achieving 53% accuracy—a 35% improvement over their original TTT method.
Building on these results, we explored various combinations of our approach with BARC’s components:
  • Combining our TTT pipeline and neural model with BARC’s synthesizer raised accuracy to 58.558.5%.
  • Combining our TTT pipeline with BARC’s neural model and synthesizer raised accuracy to 61.87561.875%.
This final configuration establishes a new state-of-the-art on the ARC public evaluation set, matching the average human performance (LeGris et al., 2024b). While this represents significant progress, there remains a substantial gap to the best human performance of 97.8%, indicating room for further improvements.

### v2

@@ 4.2 Experimental Details
For our ablation experiments, we use the 1B-parameter Llama-3.2 model (Llama Team, 2024). For our final results in Section 4.6, we use the 8B Llama 3 model. We use Low-Rank Adaptation  (LoRA; Hu et al., 2022) for parameter-efficient test-time training. More details are given in Section C.2.
Fine-tuning before TTT
While TTT offers task-specific adaptation, the initial capabilities of the base model significantly influence its final performance (Section 4.4). We developed several approaches for generating synthetic training data to enhance the base model’s abstract reasoning capabilities through fine-tuning, exploring both automated and semi-automated methods for task generation. This is complementary to TTT as the base model is fine-tuned on tasks distinct from those tested on, when TTT is applied. Details on our data generation strategies, as well as the effects of various data sources and model sizes on performance, are provided in Appendix B. The fine-tuned base model serves as the foundation for all subsequent experiments.
Evaluation
@@
The success criterion requires producing an exact match for all test outputs (no partial credit). Following the standard ARC scoring criteria, we use the pass@2 metric and produce $2$ attempts for each test input. The original training and validation sets consist of $400$ tasks each. However, for efficient evaluation purposes, we randomly pick $80$ balanced ARC tasks from the ARC validation set, including $20$ easy, $20$ medium, $20$ hard, $20$ expert tasks according to the classification in (LeGris et al., 2024) (see Table 2 for this task list). Except for our final results, we use this subset of ARC tasks throughout our experiments. We limit $\mathcal{D}_{\textrm{TTT}}$ to have a maximum of $250$ examples per task for efficiency reasons. Section C.2 provides additional details on the hyperparameters.
Inference
One of the most common techniques to scale inference-time compute is to use temperature sampling to obtain multiple responses and select the best according to a ranker, called self-consistency (Wang et al., 2023). However, this is not viable in ARC (where the output grid is directly predicted) as there is no way to directly enforce diversity across samples while ensuring coherence within samples. As an alternative self-consistency approach, we try an augmented inference strategy that combines greedy decoding with multiple versions of the input. Specifically, we generate multiple prediction candidates by using geometric transformations. We then employ a hierarchical voting strategy to determine the final prediction from the set of generated candidates. This approach involves two stages of voting to progressively narrow down the best candidates: (1) Intra-transformation voting: Group predictions by their corresponding transformation $t$. Within each group, select the top-3 most frequent predictions. (2) Global voting: Take the selected transformation-specific candidates from the previous step and select the top-2 most frequent predictions across all transformations. The augmented inference pipeline is summarized in Figure 4 and full details of the pipeline are in Appendix E.
@@
Figure 4: Augmented inference and hierarchical voting. We use leave-one-out tasks and invertible geometric transformations to obtain multiple equivalent versions of the task for augmented inference. Predictions from these versions are aggregated with a hierarchical voting strategy: first, voting is performed within each transformation, and then the top candidates from each transformation undergo global voting to yield the top two predictions.
Figure 5: Accuracy of different data and optimization ablations in TTT on ARC. Our data ablations reveal that the ICL data format is crucial for effective TTT, and that applying transformations to augment the TTT dataset notably enhances performance. For optimization, learning task-specific adapters significantly outperforms using a single adapter and taking a loss on the in-context demonstrations provides a minor performance boost.
4.3 Impact of TTT Design
In this section, we compare the final implementation of our method with different design choices for TTT. FT serves as the baseline, using only the fine-tuned model with demonstrations in-context. No Transformations omits the augmentation step. Direct I/O Data replaces in-context tasks with the direct input-output task formulation (Section 3.1). Shared TTT uses a single LoRA adapter across all tasks instead of learning one per task. No Demonstration Loss removes the loss on demonstration outputs (Section 3.1).
Results are presented in Figure 5. Our TTT method is effective, improving fine-tuned model accuracy approximately 6$\times$ $\bf(5\%\to 29\%)$. In-context formatting is especially important; using the direct input-output data to construct $\mathcal{D}_{\textrm{TTT}}$ causes an $11$-task drop ($38\%$). Removing transformations causes a $16$ task drop ($55\%$). Regarding optimization, per-task LoRA adapters outperform a single shared adapter by $7$ tasks ($24\%$). Including losses on the demonstration outputs yields a modest but consistent gain ($26\%\to 29\%$).
Figure 6: Performance results across model sizes. Fine-tuned model performance improves with increasing size. However, the scaling behavior after TTT is less clear. For instance, the final performance of the 1B and 3B models is identical after TTT.
4.4 Impact of Model Size
@@
We perform full fine-tuning of 1B and 3B Llama 3.2 (instruction-tuned) and 8B Llama 3 (instruction-tuned) using synthetically generated data, as detailed in Appendix B, and then use our default TTT implementation. We show results using different model sizes in Figure 6. Increasing the model size consistently improves FT performance, with the 8B model achieving the highest accuracy of $36\%$. At all model sizes, TTT leads to significant improvements in performance. We also observe that for smaller model sizes, TTT effectively closes the performance gap, with the 1B and 3B models achieving similar accuracy after TTT.
4.5 Impact of Augmented Inference
To analyze the impact of augmented inference and voting, we run several ablations: (1) Vanilla, which generates two predictions without transformations or advanced voting; (2) Transformed Inference, applying a single transformation (Rotate, Transpose, or Flip) to measure its isolated effect; (3) Hierarchical Voting, our full pipeline combining augmented inference and structured voting; (4) Flattened Voting, which selects the top-2 predictions from a single voting round over all generated outputs; and (5) Oracle, an upper bound that selects the correct answer if present.
As shown, individual transformations are modestly effective on their own (with Transpose performing worst), but their aggregation improves results markedly. Hierarchical voting further outperforms a flattened voting approach and closely approaches the oracle’s accuracy, suggesting that our two-stage aggregation effectively identifies the correct solution when it is present.
Figure 7: Accuracy of different transformations and voting schemes. While individual transformations generally perform at a modest level and are comparable to one another, aggregating across them through voting yields substantial improvements. Notably, a hierarchical voting strategy with two voting stages surpasses a flat voting approach. Our hierarchical method approaches oracle-level performance, demonstrating its effectiveness in accurately selecting the correct answer when present.
4.6 Comparison to Other Systems
@@
Following our experiments on $80$ tasks, we present comprehensive results on the full ARC public evaluation set, comparing our system against existing approaches. Our analysis focuses on three key aspects: the impact of our TTT methodology, the benefits of combining our approach with existing methods, and the differences between fully neural and program synthesis methods.
TTT
@@
We apply TTT and augmented inference procedure to our base fine-tuned model (fine-tuned 8B model). TTT significantly improves accuracy from $18.3$% to $47.125$%.
@@
PS Fine-tuned LM TTT Method Score
X Ours X $18.25$%
X Ours Ours $47.125$%
X BARC Ours $53$%
BARC Ours Ours $58.5$%
BARC BARC Ours $62.8$%
Avg. Human $60.2$%
Best Human $97.8$%
BARC (ensemble) $54.375$%
BARC (no synthesizer) $39.25$%
Claude 3.5 Sonnet $21$%
GPT-4o $9$%
OpenAI o1 preview $21.0$%
DeepSeek r1 $20.5$%
OpenAI o3 $82.8$%
Table 1: Pass@2 Scores of different systems on the ARC validation set. Our TTT pipeline improves base models consistently. We achieve $47.1\%$ accuracy when applied to our fine-tuned model and $53.0\%$ when applied to the BARC model (Li et al., 2025). We ensemble our method with program synthesis (PS) based models, where we achieve score of $61.875$%, comparable to the average human performance of $60.2\%$.
A concurrent work by Li et al. (2025) introduced BARC, achieving $54.375$% accuracy by combining neural and program synthesis approaches. While their fully neural approach shares similarities with our system, our TTT and inference pipeline has several additional components (per-task LoRA, more augmentations, hierarchical voting) that boost performance. To validate our improvements, we applied our TTT pipeline to BARC’s fully neural model, achieving $53.0\%$ accuracy—a $35\%$ improvement over their original TTT method.
Building on these results, we explored combinations of our approach with BARC. Combining our TTT pipeline and neural model with BARC’s synthesizer raised accuracy to $58.5$%. Combining our TTT pipeline with BARC’s neural model and synthesizer raised accuracy to $61.875$%. This configuration matches average human performance of $60.2\%$ (LeGris et al., 2024) on the benchmark.